# Mosaic3D: Foundation Dataset and Model for Open-Vocabulary 3D Segmentation

### v1

@@ 3.2 Data Statistics
  • Scale: We generate over 5.6M mask-text pairs with 30M text tokens across 30K scenes, which is significantly larger than previous datasets in terms of both scene coverage and annotation density.
@@
  • Richness: The region-aware VLM generates detailed, contextual descriptions that capture both visual attributes and spatial relationships, providing richer semantic information than simple object labels.
@@
4 Model Training
Building upon the Mosaic3D-5.6M dataset, we develop a model for open-vocabulary 3D scene understanding that performs both semantic and instance segmentation without being constrained to predefined categories. Our training follows a two-stage approach: first, we train Mosaic3D, a 3D encoder to align per-point features with text embeddings (Sec. 4.1), followed by a lightweight mask decoder that learns segment masks from these aligned features (Sec. 4.2).
@@ 4.1 Mosaic3D: Language-Aligned 3D Encoder
Architecture. As a backbone of our model, we use U-shaped sparse convnets [31, 19] which efficiently process sparse 3D data through sparse convolutions. Given a point cloud $\mathbf{P}\in\mathbb{R}^{N\times 3}$, our 3D backbone $f(\cdot)$ computes per-point feature vectors as $f(\mathbf{P})=\mathbf{Z}\in\mathbb{R}^{N\times D}$.
Training objective. To align the geometric embeddings with language semantics, we employ a contrastive learning framework, inspired by previous works [92, 25]. Given mask-text pairs $\{(\mathbf{s}_{k},c_{k})\}_{k=1}^{K}$, a pre-trained text encoder $g(\cdot)$ computes text embeddings for each caption $\mathbf{w}_{k}\in\mathbb{R}^{D}$:
$\mathbf{Z}=f(\mathbf{P}),\quad\mathbf{w}_{k}=g(c_{k}).$ (3)
The model calculates similarity scores between point embeddings and text embeddings, and pools these scores using the region masks to obtain region-level similarity scores. The final training objective consists of:
$\mathcal{L}_{\mathrm{point}}$ $=-\frac{1}{K}\sum_{k=1}^{K}\mathrm{Pool}\left(\mathbf{s}_{k},\sigma(\mathbf{Z},\mathbf{w}_{k})\right),$ (4)
$\sigma(\mathbf{Z},\mathbf{w}_{k})$ $=\log\frac{\exp(\mathbf{Z}\cdot\mathbf{w}_{k}^{\top}/\tau)}{\sum_{j=1}^{K}\exp(\mathbf{Z}\cdot\mathbf{w}_{j}^{\top}/\tau)},$
where $\operatorname{Pool(\cdot,\cdot)}$ denotes masked pooling operator and $\tau$ is a learnable logit temperature.
@@ 4.2 Mosaic3D with Mask Decoder
We train a lightweight mask decoder on top of our language-aligned backbone to enable open-vocabulary 3D instance segmentation, avoiding the need for additional instance segmentation networks used in prior work [81, 60, 39, 94].
@@
Architecture. Our mask decoder uses Mask3D [76], a Transformer-based architecture adapted from 2D segmentation [15, 16] to 3D. Specifically, our decoder takes non-parametric queries (i.e., positional encodings of points sampled from the input point cloud) and language-aligned point features from our backbone (Sec. 4.1) as inputs. The decoder outputs mask embeddings aligned with language features, enabling open-vocabulary segmentation of 3D scenes.
@@
Training data. To enable open-vocabulary 3D instance segmentation, we need training data that is not constrained to pre-defined categories. While prior works [81, 60, 39] used closed-vocabulary labels, we leverage Segment3D [37]’s class-agnostic masks predicted by SAM [47] and combine them with our multi-view mask-caption data to create a rich open-vocabulary training set. As detailed in Algorithm 1, we merge our mask-caption data $\{(\mathbf{s}_{k},c_{k})\}_{k=1}^{K}$ with Segment3D [37] masks $\{\mathbf{s}_{l}^{\mathrm{3D}}\}_{l=1}^{L}$ based on IoU matching. This yields a set of Segment3D masks $\{(\mathbf{s}^{\mathrm{3D}}_{m},\{c_{j\in\mathcal{M}_{m}}\})\}_{m=1}^{M}$, each associated with multiple captions from our multi-view data. During training, we randomly sample a fixed number of captions for each mask from its associated caption set.
@@
Training objective. Given a point cloud $\mathbf{P}$ and Segment3D masks with associated captions $\{(\mathbf{s}^{\mathrm{3D}}_{m},\{c_{j\in\mathcal{M}_{m}}\})\}_{m=1}^{M}$, we compute $Q$ numbers of mask embeddings $\mathbf{Z}_{\mathrm{mask}}\in\mathbb{R}^{Q\times D}$ and normalized text embeddings $\bar{\mathbf{w}}_{m}\in\mathbb{R}^{D}$:
$\mathbf{Z}_{\mathrm{mask}}=h(\mathbf{Z};\mathbf{Q}),\quad\mathbf{\bar{w}}_{m}=g(\bar{c}_{m}),$ (5)
@@
where $h(\cdot)$ is our mask decoder that takes point features $\mathbf{Z}=f(\mathbf{P})$ and sampled queries $\mathbf{Q}$ as input, $g(\cdot)$ is a pre-trained text encoder, and $\bar{c}_{m}$ concatenates all captions associated with mask $\mathbf{s}^{\mathrm{3D}}_{m}$. Following Segment3D [37], we first train the mask decoder to predict binary instance masks using three standard losses: objectness prediction loss $\mathcal{L}_{\mathrm{obj}}$, Dice loss $\mathcal{L}_{\mathrm{dice}}$, and binary cross entropy loss $\mathcal{L}_{\mathrm{bce}}$. After Hungarian matching, these losses are computed on:
@@
$\mathbf{o}=\mathrm{Linear}(\mathbf{Z}_{\mathrm{mask}}),\quad\mathbf{S}_{\mathrm{pred}}=\sigma(\mathbf{Z}_{\mathrm{mask}}\cdot\mathbf{Z}^{\top}),$ (6)
@@
where $\mathbf{o}\in\mathbb{R}^{Q\times 2}$ and $\mathbf{S}_{\mathrm{pred}}\in\mathbb{R}^{Q\times N}$ are objectness scores and predicted binary masks. To enable open-vocabulary segmentation, we introduce a mask caption loss $\mathcal{L}_{\mathrm{cap}}$ that explicitly aligns mask embeddings with caption embeddings:
$\mathcal{L}_{\mathrm{cap}}$ $=-\frac{1}{M}\sum_{m=1}^{M}\sigma(\mathbf{Z}_{\mathrm{mask}},\mathbf{\bar{w}}_{m}),$ (7)
$\sigma(\mathbf{Z}_{\mathrm{mask}},\mathbf{\bar{w}}_{m})$ $=\log\frac{\exp(\mathbf{Z}_{\mathrm{mask}}\cdot\mathbf{\bar{w}}_{m}^{\top}/\tau)}{\sum_{j=1}^{M}\exp(\mathbf{Z}_{\mathrm{mask}}\cdot\mathbf{\bar{w}}_{j}^{\top}/\tau)}.$
@@ 5 Experiments
In this section, we investigate how dataset size impacts model performance (Sec. 5.2). Next, we benchmark Mosaic3D against existing methods for open-vocabulary 3D scene segmentation (Sec. 5.3). Finally, we evaluate each component of the Mosaic3D data engine to analyze their contributions to the overall system performance (Sec. 5.4).
@@ 5.3 Benchmark Results
Figure 7: Attention visualization of Mosaic3D as a 3D foundational model. The attention maps between point features and text queries highlight semantically relevant regions, demonstrating Mosaic3D’s effectiveness for open-vocabulary 3D understanding.
Open-vocabulary 3D instance segmentation. As shown in Tab. 2, methods using both 2D and 3D inputs achieve strong results by directly applying CLIP models, but are impractical due to high latency (33-285 sec per scene) from processing multiple view images. Therefore, we focus our comparison on methods using only 3D inputs. For fair comparison, we evaluate our Mosaic3D trained only on ScanNet against prior methods [39, 81, 94, 60, 63, 92] that also train and evaluate on ScanNet. With Mask3D [76] as a closed-vocabulary region proposal network, Mosaic3D outperforms the previous best method OpenIns3D [39] by 3.0p mAP. When using truly open-vocabulary Segment3D [37] proposals, Mosaic3D maintains strong performance at 2.7 mAP despite the more challenging setting. Finally, our lightweight mask decoder trained with Mosaic3D-5.6M with caption merging (Alg. 1) achieves 3.9 mAP while being the first single-stage open-vocabulary 3D instance segmentation model that does not require ground truth labels.
5.4 Analysis of Data Engine Components
Finally, we conduct an ablation study on different component combinations of Mosaic3D data engine to measure the contribution of each component. To understand the impact of different components in our data generation pipeline, we conduct ablation experiments by systematically varying key components while keeping the model architecture (SparseUNet16 [92]) and training data (ScanNet only) fixed. As shown in Table 3, we evaluate the following configurations:
@@ 5.4 Analysis of Data Engine Components
Based on these experiments, we adopt Grounded-SAM [71] with SAM2 [70] and SEEM [104] for mask generation combined with Osprey [97] for captioning as our final pipeline, as this configuration achieves the best overall performance.
@@ 6 Conclusion
Limitation. While Mosaic3D-5.6M is a large-scale and fine-grained dataset for open-vocabulary 3D segmentation, its effectiveness diminishes when 2D views captured or rendered from 3D scenes become too sparse. As shown in Fig. 6, datasets with sparse views per scene like Matterport3D [12] and Structured3D [101] (Average 88 and 9 views per scene, respectively) show less improvement then densely captured datasets like ARKitScenes [5] and ScanNet++ [93]. Addressing this limitation through generative models to synthesize additional views could be an interesting future direction.
[…]
@@ C.2 Impact of Text Encoders
Among all variants, Recap-CLIP achieves the best overall performance with 68.1% f-mIoU on ScanNet20 and 15.7% f-mIoU on ScanNet200. This represents a +0.3% and +0.9% improvement over the base CLIP/B16 model respectively. The superior performance of Recap-CLIP aligns with its enhanced text-image alignment ability demonstrated in 2D vision tasks. Based on these comprehensive experiments, we select Recap-CLIP as our default text encoder for all subsequent experiments. To ensure fair comparisons with previous work, we maintain consistency by using the same text encoder configuration when reproducing baseline results, as shown in Tables 13, and A6. This standardization enables direct performance comparisons and validates the improvements achieved by our proposed approach.

### v2

@@ 3.2 Data Statistics
  • Scale: We generate over 5.6M mask-text pairs with 30M text tokens across 30K scenes, significantly larger than previous datasets in scene coverage and annotation density.
@@
  • Richness: The region-aware VLM generates detailed contextual descriptions that capture both visual attributes and spatial relationships, providing richer semantic information than simple object labels.
@@
4 Mosaic3D Model Training
We train the 3D open-vocabulary segmentation model based on Mosaic3D-5.6M using two-stage training: per-point language alignment (Sec. 4.1), and mask decoder training that predicts instances from these aligned features (Sec. 4.2).
@@ 4.1 Mosaic3D: Language-Aligned 3D Encoder
Architecture. We use U-shaped sparse convnets [33, 21] for our backbone due to their efficiency and scalability. Given a point cloud $\mathbf{P}=\{\mathbf{p}_{i}\}_{i=1}^{N},\mathbf{p}_{i}\in\mathbb{R}^{3}$, the network $\mathcal{E}^{\text{3D}}(\cdot)$ outputs per-point features $\mathcal{E}^{\text{3D}}(\mathbf{P})=\{\mathbf{z}^{\text{3D}}_{i}\}_{i=1}^{N},\mathbf{z}^{\text{3D}}_{i}\in\mathbb{R}^{D}$.
Training objective. To align the geometric embeddings with language semantics, we employ a contrastive learning framework [98, 27]. Given 3D mask-text pairs $\{(\mathbf{s}_{k},c_{k})\}_{k=1}^{K}$, a pre-trained text encoder $\mathcal{E}^{\text{text}}(\cdot)$ computes text embeddings $\mathbf{z}^{\text{text}}_{k}\in\mathbb{R}^{D}$ for each caption. The similarity scores between point features and text embeddings are averaged using the region masks to weigh all regions equally. The final training objective is:
$\mathcal{L}_{\mathrm{point}}$ $=-\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{N}(\mathbf{s}_{k})_{i}\log\frac{\exp(\mathbf{z}^{\text{3D}}_{i}\cdot\mathbf{z}^{\text{text}}_{k}/\tau)}{\sum_{j=1}^{K}\exp(\mathbf{z}^{\text{3D}}_{i}\cdot\mathbf{z}^{\text{text}}_{j})}$ (2)
where $\tau$ is a learnable logit temperature.
@@ 4.2 Mosaic3D with Mask Decoder
On top of our language-aligned backbone, we add a lightweight mask decoder to enable open-vocabulary 3D instance segmentation, avoiding the need for separate instance segmentation networks used in prior work [87, 65, 42, 100].
@@
Architecture. We use Mask3D [82] as our mask decoder, a transformer-based architecture adapted from 2D segmentation [17, 18]. Specifically, our decoder takes non-parametric queries (i.e., positional encodings of points sampled from the input point cloud) and language-aligned point features from our backbone (Sec. 4.1) as input. The decoder outputs mask embeddings aligned with language features, enabling open-vocabulary segmentation of 3D scenes.
@@
Training data. To enable open-vocabulary 3D instance segmentation, we need training data that is not constrained to predefined categories. While prior work [87, 65, 42] used closed-vocabulary labels, we leverage Segment3D [40]’s class-agnostic masks predicted by SAM [50] and combine them with our multi-view mask-caption data to create a rich open-vocabulary training set. As detailed in Algorithm 1, we merge our mask-caption data $\{(\mathbf{s}_{k},c_{k})\}_{k=1}^{K}$ with Segment3D [40] masks $\{\mathbf{s}_{l}^{\mathrm{3D}}\}_{l=1}^{L}$ based on IoU matching. This yields a set of Segment3D masks $\{(\mathbf{s}^{\mathrm{3D}}_{m},\{c_{j\in\mathcal{M}_{m}}\})\}_{m=1}^{M}$, each associated with multiple captions from our multi-view data. During training, we randomly sample a fixed number of captions for each mask from its associated caption set.
@@
Training objective. Given a point cloud $\mathbf{P}$ and Segment3D masks with associated captions $\{(\mathbf{s}^{\mathrm{3D}}_{m},\{c_{j\in\mathcal{M}_{m}}\})\}_{m=1}^{M}$, we compute $Q$ numbers of mask embeddings $\mathbf{Z}^{\mathrm{mask}}\in\mathbb{R}^{Q\times D}$ and normalized text embeddings $\bar{\mathbf{z}}^{\text{text}}_{m}\in\mathbb{R}^{D}$:
$\mathbf{Z}^{\mathrm{mask}}=\{\mathbf{z}^{\text{mask}}_{q}\}=\mathcal{D}(\{\mathbf{z}^{\text{3D}}_{i}\};\mathbf{Q}),\quad\mathbf{\bar{z}}^{\text{text}}_{m}=\mathcal{E}^{\text{text}}(\bar{c}_{m}),$ (3)
@@
where $\mathcal{D}(\cdot)$ is our mask decoder that takes point features and sampled queries $\mathbf{Q}$ as input, and $\bar{c}_{m}$ concatenates all captions associated with mask $\mathbf{s}^{\mathrm{3D}}_{m}$. Following Segment3D [40], we first train the mask decoder to predict binary instance masks using three standard losses: objectness prediction loss $\mathcal{L}_{\mathrm{obj}}$, Dice loss $\mathcal{L}_{\mathrm{dice}}$ [62], and binary cross entropy loss $\mathcal{L}_{\mathrm{bce}}$. Then, we use Hungarian matching to find the set of mask predictions that minimizes error given ground-truth masks. The losses are computed as:
@@
$\mathbf{o}=\mathrm{Linear}(\mathbf{Z}^{\mathrm{mask}}),\quad\mathbf{S}=\sigma(\mathbf{Z}^{\mathrm{mask}}\cdot\mathbf{Z}^{\text{3D}\top}),$ (4)
@@
where $\mathbf{o}\in\mathbb{R}^{Q\times 2}$ and $\mathbf{S}\in\mathbb{R}^{Q\times N}$ are objectness scores and predicted binary masks. To enable open-vocabulary segmentation, we introduce a mask caption loss $\mathcal{L}_{\mathrm{cap}}$ that explicitly aligns mask embeddings with caption embeddings:
$\mathcal{L}_{\mathrm{cap}}=-\frac{1}{M}\sum_{m=1}^{M}\log\frac{\exp(\mathbf{z}^{\text{mask}}_{m}\cdot\bar{\mathbf{z}}^{\text{text}}_{k}/\tau)}{\sum_{j=1}^{M}\exp(\mathbf{z}^{\text{mask}}_{m}\cdot\bar{\mathbf{z}}^{\text{text}}_{j})}$ (5)
@@ 5 Experiments
In this section, we investigate how dataset size impacts model performance (Sec. 5.2). Next, we benchmark Mosaic3D against existing methods for open-vocabulary 3D scene segmentation (Sec. 5.3). Finally, we analyze our system through attention visualization, zero-shot experiments, and ablation studies on the Mosaic3D-5.6M data engine (Sec. 5.4).
@@ 5.3 Benchmark Results
Open-vocabulary 3D instance segmentation. As shown in Table 2, methods using both 2D and 3D inputs achieve strong results by directly applying CLIP models, but are impractical due to high latency (33-285 sec per scene) from processing multiple view images. For fair comparison, we evaluate our Mosaic3D trained only on ScanNet against prior methods [42, 87, 100, 65, 68, 98] that also train and evaluate on ScanNet. With Mask3D [82] as a closed-vocabulary region proposal network, Mosaic3D outperforms the previous best method OpenIns3D [42] by 3.0p mAP. When using truly open-vocabulary Segment3D [40] proposals, Mosaic3D maintains strong performance at 2.7 mAP despite the more challenging setting. Finally, our lightweight mask decoder trained with Mosaic3D-5.6M with caption merging (Alg. 1) achieves 3.9 mAP while being the first single-stage open-vocabulary 3D instance segmentation model that does not require ground truth labels.
5.4 Analysis
@@
Figure 7: Attention visualization of Mosaic3D as a 3D foundational model. From left to right, the first two examples show results on ScanNet [24], while the next two examples show results on ETH3D [80]. More examples are provided in the supplmentary material.
Data engine components. Finally, we conduct an ablation study on different component combinations of Mosaic3D data engine to measure the contribution of each component. To understand the impact of different components in our data generation pipeline, we conduct ablation experiments by systematically varying key components while keeping the model architecture (SparseUNet16 [98]) and source dataset (ScanNet only) fixed. As shown in Table 3, we evaluate the following configurations:
@@
Our final pipeline combines Grounded-SAM2 [76, 75] with SEEM [110] for mask generation and Osprey [103] for captioning, as this configuration yields optimal performance.
Attention visualization. In Fig. 7, we visualize the similarity between dense point features obtained from Mosaic3D and free-form text queries. As shown in the figure, despite being trained only on indoor scenes, Mosaic3D effectively localizes semantic regions described by text queries across both indoor [24] and outdoor [80] datasets.
Zero-shot 3D semantic segmentation. While annotation-free methods [98, 48] do not utilize GT annotations, their generated captions often contain class names from the evaluation set (e.g., "chair"), making them not truly "zero-shot." To address this, we conduct a more rigorous zero-shot analysis by anonymizing all class names in the training captions, replacing them with general terms like "object." To ensure fair comparison, we kept all experimental settings (e.g., model architecture, CLIP model, loss function) identical across methods, varying only the training data. As shown in Table 4, all models experience performance drops when trained on anonymized data, but the model trained on Mosaic3D-5.6M maintains the strongest performance (10.5 f-mIoU), outperforming both annotation-free methods and approaches that use ground truth annotations like LEO [39].
Dataset f-mIoU $\Delta$
w/o Anon. w/ Anon.
RegionPLC [98] 8.5 3.8 $\downarrow$   4.7
OV3D [48] 9.2 4.3 $\downarrow$   4.9
Mosaic3D-SN [^1] A subset of Mosaic3D-5.6M using only ScanNet as source dataset 13.0 8.7 $\downarrow$   4.3
LEO [39] 14.8 2.9 $\downarrow$ 11.9
SceneVerse [44] 13.6 4.7 $\downarrow$   8.9
EmbodiedScan [94] 6.7 0.0 $\downarrow$   6.7
MMScan [60] 11.7 8.5 $\downarrow$   3.2
Mosaic3D-5.6M 15.7 10.5 $\downarrow$   5.2
Table 4: Zero-shot semantic segmentation on ScanNet200 [78]. “Anon.” refers to class name anonymization, and $\flat$ indicates our re-implementation of OV3D as the original data is not available.
@@ 6 Conclusion
Acknowledgement. This work was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants (RS-2021-II212068: AI Innovation Hub, RS-2024-00457882: National AI Research Lab Project) funded by the Korea government (MSIT).
@@ A.1 Data Statistics
In Tab. A2, we evaluate caption and 3D mask quality across datasets using three metrics. The unique normalized nouns count measures the total number of unique normalized nouns in captions, with higher count indicating richer and more diverse captions. Mask coverage (%) calculates the mean percentage of 3D points with associated captions per scene, where higher coverage enables more effective training. Mask entropy (bits) measures mask quality for datasets with partial masks generated from multi-view images (i.e. OV3D, RegionPLC, and Mosaic3D-5.6M) without using GT. It calculates Shannon entropy of GT instance ID distributions within each mask–higher entropy indicates that a mask contains multiple GT instances, suggesting less accurate mask boundaries. Mosaic3D-5.6M demonstrates superior caption diversity and mask quality compared to both existing large-scale 3D-text datasets and previous open-vocabulary 3D segmentation datasets, validating its value as a new dataset.
Train dataset Used Mask-Caption Quality ScanNet20 ScanNet200
ScanNet GT # Nouns Coverage Entropy f-mIoU f-mIoU Head Com. Tail
Datasets using only ScanNet as source
OV3D 2.5K 70.6 72.8 45.6 7.0 18.6 2.1 0.1
RegionPLC 1.4K 77.3 81.0 50.4 8.5 21.1 3.6 0.7
Mosaic3D-SN [^2] A subset of Mosaic3D-5.6M using only ScanNet as source dataset. 9.0K 92.6 60.7 65.0 13.0 30.2 6.9 1.4
Datasets using multiple sources
LEO 2.6K 66.2 - 65.9 14.8 34.3 8.3 1.4
SceneVerse 8.8K 60.0 - 67.3 13.6 32.4 7.3 0.8
EmbodiedScan 0.3K 14.0 - 44.8 6.7 16.1 3.6 0.2
MMScan 6.0K 48.0 - 64.1 11.7 26.1 7.9 0.7
Mosaic3D-5.6M 29.9K 93.7 - 68.1 15.7 32.9 10.8 2.7
Table A2: Dataset comparison. We analyze mask-caption quality metrics and annotation-free 3D semantic segmentation performance of different training datasets, while keeping the same model architecture (SpUNet-34C), CLIP model (Recap-CLIP), and loss function (Contrastive).
@@ C.2 Impact of Text Encoders
Among all variants, Recap-CLIP achieves the best overall performance with 68.1% f-mIoU on ScanNet20 and 15.7% f-mIoU on ScanNet200. This represents a +0.3% and +0.9% improvement over the base CLIP/B16 model respectively. The superior performance of Recap-CLIP aligns with its enhanced text-image alignment ability demonstrated in 2D vision tasks. Based on these comprehensive experiments, we select Recap-CLIP as our default text encoder for all subsequent experiments. To ensure fair comparisons with previous work, we maintain consistency by using the same text encoder configuration when reproducing baseline results, as shown in Tables 13, and A7. This standardization enables direct performance comparisons and validates the improvements achieved by our proposed approach.